# Mitigating Nonlinear Systematics in Weak Lensing Surveys II: Stability and Diagnostics with Intrinsic Alignment

### v1

@@ V Discussion and Conclusions
In the first scenario, the “reasonably accurate” IA case corresponds to using the same IA model for both the sampling and fiducial analyses, while certain fixed parameters of the fiducial model cannot be recovered by the sampling model. We find that the BNT approach still succeeds in separating the physical scales in $k$-space and yields an unbiased estimate of $S_{8}$.
In the second scenario, the fiducial and sampling IA models differ fundamentally—either in redshift or in $k$-space—and can therefore introduce significant biases in the cosmological constraints. The key question is whether we can diagnose inconsistencies in the IA model and thus identify potential biases in the inferred cosmological parameters. Our results show that when the two IA models differ in redshift space, both the noBNT and BNT analyses provide an internal self-consistency check that reveals the discrepancy between the fiducial and sampling models. However, when the IA models differ in $k$-space, only the BNT approach provides the correct diagnostic, the noBNT approach does not provide any means of identifying an incorrect IA model. This is indeed expected as BNT is built to offer a way to separate physical scales in weak lensing.
The main conclusion of this work is that intrinsic alignments do not cause the BNT method to fail. The separation of $k$-scales based on linear combinations of tomographic bins still effectively removes high-$k$ modes where the modelling is unreliable. We also explored the use of $k$-bins as an extension of BNT to isolate physical scales around specific features of interest. Our results show that a reasonably accurate IA model is required but need not be exact, and that when the IA model is fundamentally incorrect, BNT nonetheless provides a consistency test for identifying such cases.

### v2

@@ V Discussion and Conclusions
[…]
The main conclusion of this work is that intrinsic alignments do not cause the BNT method to fail. The separation of $k$-scales based on linear combinations of tomographic bins still effectively removes high-$k$ modes where the modelling is unreliable. We also explored the use of $k$-bins as an extension of BNT to isolate physical scales around specific features of interest. Our results show that a reasonably accurate IA model is required but need not be exact, and that when the IA model is fundamentally incorrect, BNT nonetheless provides a consistency test for identifying such cases. The robustness against biased IA models can be understood from Figures 7 and 9: in the BNT cases, the binning in $k$ or $z$ space allows the bias introduced by an IA model to be locally compensated within a given $k$ or $z$ bin by adjusting $A_{\rm TA,eff}$, without compromising the cosmological diagnostic power of the BNT approach.
A natural extension of this work would be to investigate whether the IA model can be characterized non-parametrically using the discrepancies observed across the redshift and wavenumber bins. One could envision an iterative procedure that progressively corrects these discrepancies in both $z$ and $k$ space. We leave this for future study, but our results suggest that an iterative framework, combining cosmological-parameter inference with $k$-cut analyses and enforcing self-consistency across redshift and $k$ bins using the BNT transform, could help further constrain the IA model while simultaneously reducing non-linear modelling uncertainties.